You’ve collected the following information from your favorite financial website. 52-Week PriceStock (Div)Div
Yld %PE
RatioClose
PriceNet
ChgHiLo36.99      13.09         Arch Coal 0.644.1      20.0    15.43    0.25    32.90       42.81         Laclede Grp 1.664.0      14.2    41.60    –0.31    194.90       151.71         IBM 3.001.6      14.7    192.37    1.91    41.86       23.44         JC Penney 0.801.9      45.8    41.77    0.35    29.92       22.82         Tootsie Roll 0.321.3      33.9    ??      –0.19     
According to your research, the growth rate in dividends for Arch Coal for the previous 10 years has been 5.3 percent.
 Required:
If investors feel this growth rate will continue, what is the required return for Arch Coal stock? 
Answer #1

Data given in the question is not in table format and not clear which value is what. I am assuming for Arch coal, Dividend is 0.64, dividend yield % 4.1 and closing price 15.43.

Required return = [Dividend*(1+growth rate)/closing price] + growth rate

Required return = [0.64*(1+0.053)/15.43] + 0.053 = [(0.64*1.053)/15.43] + 0.053 = (0.67392/15.43) + 0.053 = 0.044 + 0.053 = 0.097 or 9.7%